# UniMapGen: A Generative Framework for Large-Scale Map Construction from Multi-modal Data

### v1

@@ Abstract
Large-scale map construction is foundational for critical applications such as autonomous driving and navigation systems. Traditional large-scale map construction approaches mainly rely on costly and inefficient special data collection vehicles and labor-intensive annotation processes. While existing satellite-based methods have demonstrated promising potential in enhancing the efficiency and coverage of map construction, they exhibit two major limitations: (1) inherent drawbacks of satellite data (e.g., occlusions, outdatedness) and (2) inefficient vectorization from perception-based methods, resulting in discontinuous and rough roads that require extensive post-processing. This paper presents a novel generative framework, UniMapGen, for large-scale map construction, offering three key innovations: (1) representing lane lines as discrete sequence and establishing an iterative strategy to generate more complete and smooth map vectors than traditional perception-based methods. (2) proposing a flexible architecture that supports multi-modal inputs, enabling dynamic selection among BEV, PV, and text prompt, to overcome the drawbacks of satellite data. (3) developing a state update strategy for global continuity and consistency of the constructed large-scale map. UniMapGen achieves state-of-the-art performance on the OpenSatMap dataset. Furthermore, UniMapGen can infer occluded roads and predict roads missing from dataset annotations. Our code will be released.
@@ 1 Introduction
Lane-level map navigation plays a vital role in both autonomous and human driving systems, especially as intelligent transportation applications expand across cities. The construction of large-scale, high-quality vector maps has thus become a critical research challenge, as such maps are essential for enabling precise localization and route planning. Traditional Standard Definition (SD) maps lack precise lane-level information, resulting in limited navigation capabilities. While High Definition (HD) maps (Liu et al. 2023b) offer comprehensive vector and attribute data, their creation demands substantial resources, particularly in terms of specialized data collection vehicles. The use of satellite image to efficiently construct large-scale high-quality, lane-level maps that fall between SD and HD, as proposed by OpenSatMap (Zhao et al. 2024), has become a trend due to its advantage of satisfying global geographical coverage.
@@
When constructing large-scale lane-level maps from satellite image, three essential challenges are inevitable (Fig. 1 right bottom): (1) Occlusion of roads (Senlet et al. 2012). The physical occlusions (e.g., by tall buildings and trees) in satellite image can cause missing or inaccurate road predictions. (2) Outdatedness of satellite image. The irregular collection of satellite image leads to time gaps between annotations and up-to-date road conditions, which hinders map accuracy. (3) Incomplete annotations. As the definition of roads varies across different annotators, there are various inconsistent and incomplete road annotations in the datasets.
@@
Existing satellite-based methods for map construction are mainly segmentation-based approaches (e.g., SegNeXt (Zhao et al. 2024)) and detection-based vectorized methods (Liao et al. 2022; Liu et al. 2023b). They fail to address the above fundamental challenges inherent in satellite image. Moreover, they overlook critical aspects of electronic map representation, which require vector formatting with global consistency. These limitations lie in three key areas: (1) Inefficient representation. Existing segmentation-based approaches require complex post-processing, while detection-based methods use fixed-point vectors, which is limited for complex long lines and wasteful for short ones. (2) Discontinuous connecting. Both segmentation and detection methods require post-processing for connecting together multiple patches, leading to discontinuous connections (Fig. 1 top). (3) Limited flexibility. Existing frameworks demonstrate limited flexibility as they support fixed modality input (e.g., Bird’s Eye View (BEV) only). This harms the framework’s ability to improve map construction using alternative inputs, such as fresher Perspective View (PV) frames or specialized road mapping prompts.
@@
To address the above critical issues, inspired by the powerful contextual learning, inference, and flexible multi-modal fusion ability of multi-modal large language models (MLLMs) (Liu et al. 2023a; Li et al. 2023a; Chen et al. 2024b), we propose UniMapGen, a MLLM-based novel framework for large-scale map construction. As shown in Fig. 1, it offers three key innovations: (1) We reformulate lane-level map construction as a token-based generative problem using a generative model backbone. This approach produces smoother vector outputs across lane lines of varying lengths. (2) We propose a state update strategy for constructing large-scale maps. The map updated at each state relies on that of the last state. (3) Our framework accepts flexible multi-modal inputs, including PV, BEV, and text prompt inputs. PV images reduce inaccurately constructed maps caused by occlusions and outdateness in satellite image. Text prompt offers flexibility for interactive map construction, eliminating the problem of incomplete annotation. Extensive experiments on OpenSatMap (Zhao et al. 2024) demonstrate the effectiveness of UniMapGen.
@@ 2 Related Work
Large-Scale Map Construction. Traditional large-scale map construction primarily relies on semantic segmentation approaches (Cheng et al. 2017; Zhou et al. 2018; Demir et al. 2018; Jiang et al. 2022; Cheng et al. 2025). Notable advances include D-LinkNet’s (Zhou et al. 2018) multi-scale dilated convolution with pretrained ResNet34 (He et al. 2016), RoadFormer’s (Jiang et al. 2022) pyramidal deformable vision transformer, and CE-RoadNet’s (Cheng et al. 2025) cascaded CNN architecture, each improving road extraction through enhanced feature learning. RoadDA (Zhang et al. 2021) further addresses domain adaptation through adversarial learning and self-training for cross-domain performance. However, segmentation-based methods face common challenges in maintaining structural continuity and struggle with lane-level map generation. Our work addresses these limitations through a generative end-to-end framework that directly produces vectorized road maps from satellite images, eliminating post-processing artifacts and enabling accurate lane-level map generation.
Vertorized Map Construction. The evolution of vectorized map construction began with HDMapNet’s (Li et al. 2022) two-stage segmentation approach, followed by VectorMapNet’s (Liu et al. 2023b) pioneering end-to-end framework using sequential point prediction. MapTR (Liao et al. 2022) introduced unified shape modeling with parallel processing, inspiring subsequent developments: MapTRv2 (Liao et al. 2024b) enhanced training through auxiliary supervision, PivotNet (Ding et al. 2023) preserved geometry using dynamic pivotal points, and P-MapNet (Jiang et al. 2024) improved robustness via masked pre-training. Recent advances include MapTracker’s (Chen et al. 2024a) temporal consistency through query propagation and SMART’s (Ye et al. 2025) integration of map priors from SD Map and satellite image. However, these detection-based methods rely on fixed-point vector representations regardless of line length, leading to potential quality degradation for very short or long lines. To address this, we propose a equidistant sampling strategy combined with a generative framework that supports variable-length line prediction, enabling more accurate and adaptive vector representations.
@@ 3 Method
This paper aims to construct large-scale lane-level maps using flexible multi-modal data, primarily relying on city-level satellite image. An ideal map can be represented as a collection of vectors that are complete and smooth, with global continuity and consistency among them (Ovsjanikov et al. 2012). However, previous methods suffer from inefficient representation, discontinuous connecting, limited flexibility, and drawbacks of satellite image (e.g., occlusion) in map construction. To solve these, we introduce UniMapGen, a generative framework to construct globally continuous and consistent maps from multi-modal data inputs (e.g., BEV, PV frames, and text) in an iterative state update strategy, overcoming challenges like occlusion and data outdateness.
@@ 3.2 Framework of UniMapGen
Since generative LLMs only support serialized data formats, as shown in Fig. 2(b), we serialize map vectors into a single sequence by: (1) resampling original vector representations, and (2) concatenating these discrete vectors in a specified order. Traditional fixed-point sampling strategies, such as the 20-point sampling used in MapTR (Liao et al. 2022), struggle to accurately represent long and complex lines in satellite images due to limited point num, while being inefficient for short lines (e.g., 2 m), where many points are redundant. To improve this, we propose an equal-distance sampling strategy that samples points at fixed distances, enabling efficient adaptive point allocation based on line length. In concatenation, we reorder map vectors for order consistency, facilitating easier learning by the LLM, and potentially simplifying map construction.
@@
A. Encoders/LLM and Tokenizers. (1) BEV Encoder. We employ ViT-Large-14 from Dinov2 (Liu et al. 2024) as our pretrained BEV encoder, encoding a single satellite image $I^{BEV}$ into N D-dimensional tokens $\boldsymbol{e}^{BEV}\in\mathbb{R}^{D\times N}$. (2) PV Encoder. We utilize a 3D convolution (Tran et al. 2015) followed by the image encoder from Qwen2-VL-2B (Wang et al. 2024) as our PV Encoder, processing L PV frames $I^{PV}$ into M D-dimensional tokens $\boldsymbol{e}^{PV}\in\mathbb{R}^{D\times M}$. (3) Vector Tokenizer and Detokenizer. We propose a special vector tokenizer to convert the previous state map $G_{n-1}$ into special tokens $\boldsymbol{e}^{G}$, subsequently processed through word embeddings. The vector detokenizer performs the inverse operation, decoding the MLLM-predicted tokens back into vectorized maps. (4) Large Language Model and Text Encoder. We adopt Qwen2.5-1.5B (Qwen et al. 2025) as our LLM, receiving multimodal input to construct maps. Additionally, we utilize the LLM’s text tokenizer to tokenize $T^{prompt}$ into $\boldsymbol{e}^{T}$. More model details are listed in suppl.
@@
MapTRsat (Liao et al. 2022) vector-based 18.20 22.13 26.36 28.25 6.02 14.12 4.18 34.10
MapTRv2sat (Liao et al. 2024b) vector-based 19.34 23.44 26.98 29.89 6.45 14.89 4.47 35.42
SegNeXt (Zhao et al. 2024) segmentation-based 20.30 25.93 29.50 31.38 6.98 16.05 5.26 33.69
@@
3.3 State Update for Global Construction
4 Experiments
Dataset. We evaluate the performance of UniMapGen on OpenSatMap dataset at 20-level, which includes 1180 training and 393 validation satellite image of 4096*4096 pixels. Each image in the dataset is annotated with lane-level vectorized maps, with each line classified into three line categories (Curb, Lane line, and Virtual Line), and eight attributes (e.g., line types and level of occlusion). To meet the large training data requirement for MLLMs, we conduct overlapped and inclined cropping augmentation, resulting in 700k training satellite patches. Details of the training data construction and augmentation are listed in the suppl..
Training Data Preparation. To support multi-mode training, beyond BEV-only mode, we prepare other training data across different modalities as follows: (1) PV-only training. We leverage the PV frames from nuSences (Caesar et al. 2020) and Argoverse2 (Wilson et al. 2023) datasets, enabling map construction from PV frames. To ensure the consistency of training data, we use the modified annotation in OpenSatMap dataset. To be specific, we first use the GPS information provided by nuSences, Argoverse2, and OpenSatMap dataset to link the PV frames to the corresponding BEV images and its annotation. Then, similar to MapTR, we crop the lines within 60m height and 30m width as ground truth, considering the perception field of PV frames. (2) PV and BEV joint training. Similarly, GPS information is also used to link PV frames to BEV images between nuSences/Argoverse2 and OpenSatMap dataset. The PV position coordinates in the BEV image and the angle representing the vehicle direction are also incorporated into the input to facilitate more accurate alignment between PV and BEV images. (3) Text-prompted target map generation. We use a semi-automatic annotation pipeline: automatic data synthesis via Segment Anything Model (SAM) (Kirillov et al. 2023), followed by human verification and correction for accurate target lane lines and text prompts matching. Additionally, we also incorporate the trace points from nuSences and Argoverse2 as text prompt for target map generation. More data preparation details and text prompts are listed in suppl.
Metric. Following (Zhao et al. 2024), the line width is set to 6 pixels by default, and evaluations are conducted at both semantic and instance levels. For semantic-level evaluation, we use the mean intersection over union (mIoU) (Everingham et al. 2010) as the metric across different line categories. For instance-level evaluation, we employ average precision (AP) metrics, which include Mask AP ($\text{AP}^{\text{M}}$) using mask IoU thresholds to determine true positive samples (Cheng et al. 2022), and Chamfer AP ($\text{AP}^{\text{C}}_{\text{D}}$(Zhao et al. 2024) by choosing the distance threshold D. Since standard AP computation requires class probabilities for each predicted line, we propose a pseudo-scoring mechanism based on the MLLM’s token prediction confidence. Specifically, for the i-th line, we define its pseudo-score $p_{i}$ as:
$p_{i}=\max(\text{softmax}(\mathbf{o}_{i}))$ (5)
4.1 Quantitative Evaluation
We compare UniMapGen with the current state-of-the-art method SegNeXt (Zhao et al. 2024) and vector-based map construction methods (e.g., MapTR) on the OpenSatMap20 dataset. SegNeXt employs a three-stage pipeline, including semantic segmentation and two post-processing processes (instance detection and vectorization). It constructs maps for 4096×4096 images by dividing them into 1024×1024 patches and predicting lines for each patch. MapTR (Liao et al. 2022) and MapTRv2 are originally PV frames-based map construction methods. To evaluate their performance of BEV-based generation, we replace the BEVFormer with ViT-Large-14 from Dinov2 (Oquab et al. 2023), followed by an MLP, to directly extract BEV features from the satellite image. Tab. 1 shows that UniMapGen outperforms SegNeXt and MapTR series across most metrics, including mIoU and APs, demonstrating the powerful ability of UniMapGen in large-scale map construction. Qualitative comparisons in Fig. 4 (a) further reveal that UniMapGen generates smoother, more continuous, and complete lines compared to SegNeXt’s post-processing and MapTR series results. Fig. 4 (g) shows an example of our constructed global map.
@@ 4.1 Quantitative Evaluation
TopoNet (Li et al. 2023b) 52.5 47.1 49.6 46.9 10.8 17.5
LaneGAP (Liao et al. 2024a) 49.9 57.0 53.2 74.1 34.9 47.5
RNTR (Lu et al. 2023) 57.1 42.7 48.9 63.7 45.2 52.8

### v2

@@ Abstract
Large-scale map construction plays a vital role in applications like autonomous driving and navigation systems. Traditional large-scale map construction approaches mainly rely on costly and inefficient special data collection vehicles and labor-intensive annotation processes. While existing satellite-based methods have demonstrated promising potential in enhancing the efficiency and coverage of map construction, they exhibit two major limitations: (1) inherent drawbacks of satellite data (e.g., occlusions, outdatedness) and (2) inefficient vectorization from perception-based methods, resulting in discontinuous and rough roads that require extensive post-processing. This paper presents a novel generative framework, UniMapGen, for large-scale map construction, offering three key innovations: (1) representing lane lines as discrete sequence and establishing an iterative strategy to generate more complete and smooth map vectors than traditional perception-based methods. (2) proposing a flexible architecture that supports multi-modal inputs, enabling dynamic selection among BEV, PV, and text prompt, to overcome the drawbacks of satellite data. (3) developing a state update strategy for global continuity and consistency of the constructed large-scale map. UniMapGen achieves state-of-the-art performance on the OpenSatMap dataset. Furthermore, UniMapGen can infer occluded roads and predict roads missing from dataset annotations. Our code will be released.
@@ 1 Introduction
Lane-level map navigation is foundational for critical applications such as autonomous driving and human driving systems. especially as intelligent transportation applications expand across cities. The construction of large-scale, high-quality vector maps has thus become a critical research challenge, as such maps are essential for enabling precise localization and route planning. Traditional Standard Definition (SD) maps lack precise lane-level information, resulting in limited navigation capabilities. While High Definition (HD) maps (liu2023vectormapnet) offer comprehensive vector and attribute data, their creation demands substantial resources, particularly in terms of specialized data collection vehicles. The use of satellite image to efficiently construct large-scale high-quality, lane-level maps that fall between SD and HD, as proposed by OpenSatMap (zhao2024opensatmap), has become a trend due to its advantage of satisfying global geographical coverage.
@@
When constructing large-scale lane-level maps from satellite image, three essential challenges are inevitable (Fig. 1 right bottom): (1) Occlusion of roads (6460256). The physical occlusions (e.g., by tall buildings and trees) in satellite image can cause missing or inaccurate road predictions. (2) Outdatedness of satellite image. The irregular collection of satellite image leads to time gaps between annotations and up-to-date road conditions, which hinders map accuracy. (3) Incomplete annotations. As the definition of roads varies across different annotators, there are various inconsistent and incomplete road annotations in the datasets.
@@
Existing satellite-based methods for map construction are mainly segmentation-based approaches (e.g., SegNeXt (zhao2024opensatmap)) and detection-based vectorized methods (liao2022maptr; liu2023vectormapnet). They fail to address the above fundamental challenges inherent in satellite image. Moreover, they overlook critical aspects of electronic map representation, which require vector formatting with global consistency. These limitations lie in three key areas: (1) Inefficient representation. Existing segmentation-based approaches require complex post-processing, while detection-based methods use fixed-point vectors, which is limited for complex long lines and wasteful for short ones. (2) Discontinuous connecting. Both segmentation and detection methods require post-processing for connecting together multiple patches, leading to discontinuous connections (Fig. 1 top). (3) Limited flexibility. Existing frameworks demonstrate limited flexibility as they support fixed modality input (e.g., Bird’s Eye View (BEV) only). This harms the framework’s ability to improve map construction using alternative inputs, such as fresher Perspective View (PV) frames or specialized road mapping prompts.
@@
To address the above critical issues, inspired by the powerful contextual learning, inference, and flexible multi-modal fusion ability of multi-modal large language models (MLLMs) (liu2023visual; li2023blip; chen2024internvl), we propose UniMapGen, a MLLM-based novel framework for large-scale map construction. As shown in Fig. 1, it offers three key innovations: (1) We reformulate lane-level map construction as a token-based generative problem using a generative model backbone. This approach produces smoother vector outputs across lane lines of varying lengths. (2) We propose a state update strategy for constructing large-scale maps. The map updated at each state relies on that of the last state. (3) Our framework accepts flexible multi-modal inputs, including PV, BEV, and text prompt inputs. PV images reduce inaccurately constructed maps caused by occlusions and outdateness in satellite image. Text prompt offers flexibility for interactive map construction, eliminating the problem of incomplete annotation. Extensive experiments on OpenSatMap (zhao2024opensatmap) demonstrate the effectiveness of UniMapGen.
@@ 2 Related Work
Large-Scale Map Construction. Traditional large-scale map construction primarily relies on semantic segmentation approaches (7873262; zhou2018d; demir2018deepglobe; jiang2022roadformer; cheng2025roadnet). Notable advances include D-LinkNet’s (zhou2018d) multi-scale dilated convolution with pretrained ResNet34 (he2016deep), RoadFormer’s (jiang2022roadformer) pyramidal deformable vision transformer, and CE-RoadNet’s (cheng2025roadnet) cascaded CNN architecture, each improving road extraction through enhanced feature learning. RoadDA (zhang2021stagewise) further addresses domain adaptation through adversarial learning and self-training for cross-domain performance. However, segmentation-based methods face common challenges in maintaining structural continuity and struggle with lane-level map generation. Our work addresses these limitations through a generative end-to-end framework that directly produces vectorized road maps from satellite images, eliminating post-processing artifacts and enabling accurate lane-level map generation.
Vertorized Map Construction. The evolution of vectorized map construction began with HDMapNet’s (li2022hdmapnet) two-stage segmentation approach, followed by VectorMapNet’s (liu2023vectormapnet) pioneering end-to-end framework using sequential point prediction. MapTR (liao2022maptr) introduced unified shape modeling with parallel processing, inspiring subsequent developments: MapTRv2 (liao2024maptrv2) enhanced training through auxiliary supervision, PivotNet (ding2023pivotnet) preserved geometry using dynamic pivotal points, and P-MapNet (jiang2024p) improved robustness via masked pre-training. Recent advances include MapTracker’s (chen2024maptracker) temporal consistency through query propagation and SMART’s (ye2025smart) integration of map priors from SD Map and satellite image. However, these detection-based methods rely on fixed-point vector representations regardless of line length, leading to potential quality degradation for very short or long lines. To address this, we propose a equidistant sampling strategy combined with a generative framework that supports variable-length line prediction, enabling more accurate and adaptive vector representations.
@@ 3 Method
This paper aims to construct large-scale lane-level maps using flexible multi-modal data, primarily relying on city-level satellite image. An ideal map can be represented as a collection of vectors that are complete and smooth, with global continuity and consistency among them (ovsjanikov2012functional). However, previous methods suffer from inefficient representation, discontinuous connecting, limited flexibility, and drawbacks of satellite image (e.g., occlusion) in map construction. To solve these, we introduce UniMapGen, a generative framework to construct globally continuous and consistent maps from multi-modal data inputs (e.g., BEV, PV frames, and text) in an iterative state update strategy, overcoming challenges like occlusion and data outdateness.
@@ 3.2 Framework of UniMapGen
Since generative LLMs only support serialized data formats, as shown in Fig. 2(b), we serialize map vectors into a single sequence by: (1) resampling original vector representations, and (2) concatenating these discrete vectors in a specified order. Traditional fixed-point sampling strategies, such as the 20-point sampling used in MapTR (liao2022maptr), struggle to accurately represent long and complex lines in satellite images due to limited point num, while being inefficient for short lines (e.g., 2 m), where many points are redundant. To improve this, we propose an equal-distance sampling strategy that samples points at fixed distances, enabling efficient adaptive point allocation based on line length. In concatenation, we reorder map vectors for order consistency, facilitating easier learning by the LLM, and potentially simplifying map construction.
@@
Format
Vector {’points’: [[257,49],…,[376,15]],’category’:’Curb’,…}
Tokens
<{><points><:><[><[><257><,><49><]>,…,<[><376>
<,><15><]><]><,><category><:><Curb>,…<}>
A. Encoders/LLM and Tokenizers. (1) BEV Encoder. We employ ViT-Large-14 from Dinov2 (remoteclip) as our pretrained BEV encoder, encoding a single satellite image $I^{BEV}$ into N D-dimensional tokens $\boldsymbol{e}^{BEV}\in\mathbb{R}^{D\times N}$. (2) PV Encoder. We utilize a 3D convolution (tran2015learning) followed by the image encoder from Qwen2-VL-2B (wang2024qwen2vlenhancingvisionlanguagemodels) as our PV Encoder, processing L PV frames $I^{PV}$ into M D-dimensional tokens $\boldsymbol{e}^{PV}\in\mathbb{R}^{D\times M}$. (3) Vector Tokenizer and Detokenizer. We propose a special vector tokenizer to convert the previous state map $G_{n-1}$ into special tokens $\boldsymbol{e}^{G}$, subsequently processed through word embeddings. The vector detokenizer performs the inverse operation, decoding the MLLM-predicted tokens back into vectorized maps. Tab. 1 illustrates an example. Specifically, each coordinate number is encoded into special tokens as a whole rather than split into single digits. For example, the number 257 will be tokenized into <257> rather than <2><5><7>. Additionally, the word will be tokenized as a whole as well, such as converting ’points’ to <points>. Finally, we remove all the spaces in the token string to reduce the number of tokens. On the contrary, the vector detokenizer performs the inverse operation, decoding the MLLM-predicted tokens (like the second line) back into vectorized maps (like the first line). (4) Large Language Model and Text Encoder. We adopt Qwen2.5-1.5B (qwen2025qwen25technicalreport) as our LLM, receiving multimodal input to construct maps. Additionally, we utilize the LLM’s text tokenizer to tokenize $T^{prompt}$ into $\boldsymbol{e}^{T}$. More model details are listed in suppl.
@@
MapTRsat (liao2022maptr) vector-based 18.20 22.13 26.36 28.25 6.02 14.12 4.18 34.10
MapTRv2sat (liao2024maptrv2) vector-based 19.34 23.44 26.98 29.89 6.45 14.89 4.47 35.42
SegNeXt (zhao2024opensatmap) segmentation-based 20.30 25.93 29.50 31.38 6.98 16.05 5.26 33.69
@@
TopoNet (li2023graph) 52.5 47.1 49.6 46.9 10.8 17.5
LaneGAP (liao2024lane) 49.9 57.0 53.2 74.1 34.9 47.5
RNTR (lu2023translating) 57.1 42.7 48.9 63.7 45.2 52.8
@@
3.3 State Update for Global Construction
4 Experiments
Dataset. We evaluate the performance of UniMapGen on OpenSatMap dataset at 20-level, which includes 1180 training and 393 validation satellite image of 4096*4096 pixels. Each image in the dataset is annotated with lane-level vectorized maps, with each line classified into three line categories (Curb, Lane line, and Virtual Line), and eight attributes (e.g., line types and level of occlusion). To meet the large training data requirement for MLLMs, we conduct overlapped and inclined cropping augmentation, resulting in 700k training satellite patches. Details of the training data construction are listed in the suppl..
Data Augmentation. As the training images of OpenSatMap are quite limited (1180 at 4096*4096). If they are cropped to 896*896 resolution without augmentation, there are only 18880 patches, which is hard to meet the large-scale training data request of Multi-modal Large Language Model. The risk of overfitting the MLLM is heightened. To mitigate this, as shown in Fig. 4, we apply an overlapped and inclined cropping data augmentation to increase the number of training patches from 2K to 700K. The i,j-th cropped patch is formatted as:
$patch_{i,j}=\text{Crop}(I^{BEV},C_{i,j},\theta),$ (5)
Training Data Preparation. To support multi-mode training, beyond BEV-only mode, we prepare other training data across different modalities as follows: (1) PV-only training. We leverage the PV frames from nuSences (caesar2020nuscenes) and Argoverse2 (wilson2023argoverse) datasets, enabling map construction from PV frames. To ensure the consistency of training data, we use the modified annotation in OpenSatMap dataset. To be specific, we first use the GPS information provided by nuSences, Argoverse2, and OpenSatMap dataset to link the PV frames to the corresponding BEV images and its annotation. Then, similar to MapTR, we crop the lines within 60m height and 30m width as ground truth, considering the perception field of PV frames. (2) PV and BEV joint training. Similarly, GPS information is also used to link PV frames to BEV images between nuSences/Argoverse2 and OpenSatMap dataset. The PV position coordinates in the BEV image and the angle representing the vehicle direction are also incorporated into the input to facilitate more accurate alignment between PV and BEV images. (3) Text-prompted target map generation. We use a semi-automatic annotation pipeline: automatic data synthesis via Segment Anything Model (SAM) (kirillov2023segment), followed by human verification and correction for accurate target lane lines and text prompts matching. Additionally, we also incorporate the trace points from nuSences and Argoverse2 as text prompt for target map generation. More data preparation details and text prompts are listed in suppl.
Metric. Following (zhao2024opensatmap), the line width is set to 6 pixels by default, and evaluations are conducted at both semantic and instance levels. For semantic-level evaluation, we use the mean intersection over union (mIoU) (everingham2010pascal) as the metric across different line categories. For instance-level evaluation, we employ average precision (AP) metrics, which include Mask AP ($\text{AP}^{\text{M}}$) using mask IoU thresholds to determine true positive samples (cheng2022masked), and Chamfer AP ($\text{AP}^{\text{C}}_{\text{D}}$(zhao2024opensatmap) by choosing the distance threshold D. Since standard AP computation requires class probabilities for each predicted line, we propose a pseudo-scoring mechanism based on the MLLM’s token prediction confidence. Specifically, for the i-th line, we define its pseudo-score $p_{i}$ as:
$p_{i}=\max(\text{softmax}(\mathbf{o}_{i}))$ (6)
4.1 Quantitative Evaluation
We compare UniMapGen with the current state-of-the-art method SegNeXt (zhao2024opensatmap) and vector-based map construction methods (e.g., MapTR) on the OpenSatMap20 dataset. SegNeXt employs a three-stage pipeline, including semantic segmentation and two post-processing processes (instance detection and vectorization). It constructs maps for 4096×4096 images by dividing them into 1024×1024 patches and predicting lines for each patch. MapTR (liao2022maptr) and MapTRv2 are originally PV frames-based map construction methods. To evaluate their performance of BEV-based generation, we replace the BEVFormer with ViT-Large-14 from Dinov2 (oquab2023dinov2), followed by an MLP, to directly extract BEV features from the satellite image. Tab. 2 shows that UniMapGen outperforms SegNeXt and MapTR series across most metrics, including mIoU and APs, demonstrating the powerful ability of UniMapGen in large-scale map construction. Qualitative comparisons in Fig. 4 (a) further reveal that UniMapGen generates smoother, more continuous, and complete lines compared to SegNeXt’s post-processing and MapTR series results. Fig. 4 (g) shows an example of our constructed global map. Additionally, Fig. 5 shows several complex examples generated by UniMapGen from 4 different cities.